OpenStudy (anonymous):

\[\int\limits {3 \over (1+2x)^2}dx\]

OpenStudy (lgbasallote):

sometimes i wonder what your grade level is...trigonometry then suddenly calculus lol =)) anyway..why not try to let u= 1 + 2x then do a u-substitution? you know how to do that right?

OpenStudy (anonymous):

Umm. Mixture of grades... Um why can't you do it like this? I tried this, but got the wrong answer.\[3 \int\limits (1+2x)^{-2}\]\[3 \int\limits (1+2x)^{-1} \times 2\]??

OpenStudy (lgbasallote):

you can...but you also have to multiply the integral by 1/2 so you maintain equalit...you multiplied by 2 to get a perfect integral..then you also pull a 1/2 outside the integral make sense?

OpenStudy (anonymous):

Why do you bring a half outside?

OpenStudy (anonymous):

let me try my first integral ... \[\LARGE \int \frac{3}{(1+2x)^2}dx=\] u=1+2x 2dx=du dx=1/2du \[\LARGE \int \frac{3}{u^2}\cdot \frac12du =\frac32\int u^{-2}du\] am I right till here? lol hahahaha....

OpenStudy (kropot72):

Answer is :\[-3/2(1+2x)^{-1}\]

OpenStudy (anonymous):

Why times by half though?

OpenStudy (lgbasallote):

u = (1 +2x) du = 2dx right? but you dont have a 2..so you multiply your integral by 2 to get a perfect integral...but this is no longer your original equation..you need to multiply by 1/2 because 1/2 times 2 is 1 so it's like nothing happened....but now you have a perfect integral in (1+2x)^-2 2dx right?

OpenStudy (anonymous):

@lgbasallote is my integral right or not?

OpenStudy (kropot72):

You need to times by half because if you don't, when you differentiate the result to check ,you have double the original expression.

OpenStudy (lgbasallote):

it's right :D

OpenStudy (anonymous):

aawww... I feel like a BOSS ahahah... lol

OpenStudy (anonymous):

OK, can you write the steps out one by one (literally again)... not sure I completely understand... Sorry!!

OpenStudy (lgbasallote):

let u = (1+ 2x) du = 2dx follow that so far?

OpenStudy (anonymous):

here's why it's multiplied by 1/2 when \(\LARGE 1+2x=u\) now we take derivative of1+2x so we have 2 then we take derivative of "u" which is 1 so 2dx=1*du 2dx=du then we multiply both sides by 2 dx=1/2 du (my version) ..

OpenStudy (lgbasallote):

your original equation was \(\LARGE \int \frac{3dx}{(1+2x)^2}\) \(\LARGE 3\int \frac{dx}{(1+2x)^2}\) i think you got the part up to here

OpenStudy (anonymous):

Yeh

OpenStudy (lgbasallote):

right now you cannot integrate that...but if your numerator had 2dx then it would be integrable by the general power formula \(\LARGE 3\int \frac{2dx}{(1+2x)^2}\) you got that right?

OpenStudy (anonymous):

Yes

OpenStudy (lgbasallote):

but that's not equal to your original equation! you added a 2!! but!! if it were.. \(\LARGE 3\int \frac{\frac{1}{2} 2dx}{(1+2x)^2}\) your numerator is just equal to 1 right? so it's as if you didnt change anything...you get that?

OpenStudy (anonymous):

Yeah

OpenStudy (lgbasallote):

but you still need 2dx so it is a perfect integral...so take out the 1/2 \(\LARGE \frac{3}{2} \int \frac{2dx}{(1+2x)^2}\) you get that?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so now you can integrate it by \(\LARGE \int (1+2x)^{-2} \) power formula... \(\LARGE \frac{(1+2x)^{-2 +1}}{-2 +1}\) then multiply it by your 3/2

OpenStudy (anonymous):

Ok. Thank you!!!

OpenStudy (dumbcow):

haha yeah its like the chain rule backwards...thats why its simpler to just use u substitution, less chance of making mistake

OpenStudy (lgbasallote):

yeah..i hate it when teachers dont just teach that :/ it's confusing when you have more values